A man spends in 3 months what he earns in 2 months. IF his savings his rs. 125 per month. Find his monthly income ?. Q2. If a certain number is multiplied by 5 and 7 is added to the product and sum is divided by 3, then 3 is subtracted from the quotient, the result is 11. Find the number ?
Answer by nerdybill(7384)

A man spends in 3 months what he earns in 2 months. IF his savings his rs. 125 per month. Find his monthly income ?
Let x = monthly (rs) income
then
total amount earned in 3 months = 3x
amount spent in 3 months = 2x
savings per month = (3x-2x)/3 = x/3
.
x/3 = 125
x = 3*125
x = rs 375
.
Q2. If a certain number is multiplied by 5 and 7 is added to the product and sum is divided by 3, then 3 is subtracted from the quotient, the result is 11. Find the number ?
Let n = the number
then
(5n+7)/3 - 3 = 11
(5n+7)/3 = 14
5n+7 = 42
5n = 35
n = 7
.